What is the largest four-digit number such that the second digit is one-fourth of the third digit, the third digit is twice the first digit, and the last digit is the same as the first digit?
Thank you!

Answer by RAY100(1637)

Let,,w,x,y,z,,,be four digits in order.
.
x=1/4 y
.
y=2w,,,,therefore x = 1/4 * 2w = w/2
.
z=w
.
rewriting the digits,,,,,w,,w/2,,,2w,,,w
.
2w is "largest" digit,,,,like to set to 9 but also have a digit, (w/2) ,,,therefore 2w =8 is largest possible.
.
restating digits,,,4,,,2,,,8,,,4,
.
checking, numbers match criteria.